# TARDis: Time Attenuated Representation Disentanglement for Incomplete Multi-Modal Tumor Segmentation and Classification

### v2

@@ 1 Introduction
Tumor segmentation in contrast-enhanced Computed Tomography (CT) is fundamentally dependent on the transient physiological dynamics of the contrast agent [bi2019artificial]. Crucially, the diagnostic boundaries of tumors are often defined by this time-dependent contrast-enhancement signal, which is unavailable in non-contrast acquisitions or incorrectly timed scans [bae2010intravenous]. Accurately capturing these dynamics usually requires a multi-phase acquisition protocol.
@@
However, obtaining a standardized, complete set of phases is rarely feasible in real-world clinical practice. First, strict limits on radiation dose and scanning time often force clinicians to skip specific phases, creating a significant missing modality problem. Second, and perhaps more challenging, is the variability of acquisition protocols across different medical centers and tumor types. The definition of a specific phase often relies on fixed time delays that do not account for patient-specific cardiac output or variations in scanner hardware. As a result, a scan labeled as arterial at one hospital may effectively be a late arterial or early venous phase at another. This temporal mismatch introduces severe domain shifts, where the intensity distribution of the same modality varies drastically between datasets. Standard deep learning frameworks, which assume fixed, synchronized input channels, degrade severely when facing such inconsistent or incomplete data [javed2024robustness].
@@
Existing approaches to multi-modal segmentation typically rely on channel-concatenation strategies, which assume a fixed set of inputs. When a modality is missing, these models often fail or require retraining. To address this, recent incomplete modality frameworks have been proposed, utilizing techniques such as arithmetic mean fusion, knowledge distillation, or shared latent space mapping [dou2020unpaired, chen2019robust]. While these methods allow for flexible inputs, they generally treat different contrast phases as independent, uncorrelated channels or simple statistical variations. By ignoring the temporal dependency between phases, existing methods struggle to infer the missing dynamic information effectively, often resulting in suboptimal feature representation [shen2019brain, sharma2019missing].
@@
In this work, we propose Time Attenuated Representation Disentanglement (TARDis), a novel framework that addresses the incomplete modality challenge by modeling the underlying hemodynamics of CT. Our central hypothesis is that the Hounsfield Unit (HU) intensity can be explicitly decoupled into a time-invariant static component and a time-dependent dynamic component. Rather than treating missing modalities as empty data channels, we treat them as missing samples on the Time-Attenuation Curve (TAC). Our method employs a shared encoder to process a flexible number of input modalities. To learn robust representations, we introduce a dual-path disentanglement mechanism, as shown in Fig. 1. First, we extract a modal-agnostic anatomical representation using a learnable embedding dictionary, ensuring consistency across different scans. Second, we model the modal-specific dynamic representation by designing a physics-aware Conditional Variational Autoencoder (CVAE) [sohn2015learning], which is conditioned on the anatomy and the estimated scan time. This allows the network to probabilistically reconstruct the dynamic enhancement features of the tumor even when specific phases are missing. By explicitly disentangling anatomy from time-dependent perfusion, TARDis learns a physically grounded latent space that is robust to missing data, superior to standard fusion or masking strategies.
@@ 2.1 Multi-Modal Medical Image Segmentation
Multi-modal learning has become a cornerstone in medical image analysis, as combining complementary information from different modalities significantly improves diagnostic precision [9363915]. Early approaches focused on input-level fusion, where registered modalities are concatenated as multi-channel inputs to a Convolutional Neural Network [kamnitsas2017efficient, milletari2016v]. While effective, this assumes a fixed set of inputs, making the model fragile to missing data. Feature-level fusion strategies have subsequently emerged to better exploit cross-modal correlations. For instance, Dolz et al. [dolz2018hyperdense] proposed HyperDenseNet to facilitate dense connectivity across multi-modal streams. More recently, Transformer-based architectures have been adopted to model long-range dependencies across modalities. Hatamizadeh et al. [hatamizadeh2022unetr] introduced UNETR, which utilizes a vision transformer encoder to fuse volumetric data. Despite their success, these complete-modality frameworks rely on the assumption that all training and testing modalities are available. In clinical practice—particularly in multi-phase CT—scans are often incomplete due to radiation dose concerns or varying acquisition protocols, rendering these rigid fusion models ineffective without retraining or zero-filling strategies that often introduce artifacts.
@@ 2.2 Learning with Incomplete Modalities
To address the missing modality problem, recent research has shifted toward flexible frameworks capable of handling arbitrary subsets of inputs. The pioneering work of HeMIS [havaei2016hemis] proposed learning a shared latent representation where first- and second-order statistics of available modalities are fused. This statistical fusion concept was extended by Dorent et al. [dorent2019hetero] in the U-HVED framework, which integrates a multi-modal VAE to reconstruct missing inputs. Other approaches employ knowledge distillation (KD) to transfer information from a complete-modality teacher to an incomplete-modality student [hu2020knowledge, wang2021acn]. However, KD methods typically require a complete set of modalities during the training phase, which may not always be available in retrospective cohorts. More recently, network architecture designs have moved toward adaptive feature fusion. RFNet [ding2021rfnet] utilizes region-aware fusion to selectively aggregate features based on image content. With the rise of Transformers, methods like mmFormer [zhang2022mmformer] and M2FTrans [m2ftrans] employ masked attention mechanisms to learn intra- and inter-modal correlations, effectively masking out missing tokens. M3AE [liu2023m3ae] adapts the Masked Autoencoder (MAE) paradigm to multi-modal settings. While these methods achieve state-of-the-art results, they generally treat modalities as independent, uncorrelated channels.
@@ 2.3 Representation Disentanglement and Generative Modeling
Representation disentanglement aims to decompose data into distinct, explanatory factors, typically separating spatial structure from appearance. In natural image synthesis, methods like MUNIT [huang2018multimodal] and DRIT++ [lee2018diverse] successfully disentangle content and style codes. In medical imaging, this paradigm has been applied to improve robustness and interpretability. Chartsias et al. [chartsias2017multimodal] utilized disentanglement for synthesizing missing MRI sequences. SDNet [chartsias2019disentangled] further decomposes medical images into spatial anatomy and non-spatial modality factors to improve segmentation generalization. VQ-VAE [van2017neural] demonstrates that quantizing latent features can effectively remove noise and preserve essential structural information. In the context of missing data, Generative Adversarial Networks and VAEs are often used to impute missing modalities before segmentation [zhou2020hi, nie2017medical]. However, most existing generative imputation methods perform image-to-image translation without explicit physical grounding.
@@ 3.3 Modal-Agnostic Representations
The static component represents the inherent anatomy, which should be consistent across all phases. However, features extracted from different contrast phases often contain phase-specific noise that obscures this underlying anatomy. We employ VQ-VAE [van2017neural] for the anatomical branch because anatomy is deterministic and time-invariant. By forcing features to map to a finite, discrete Embedding Dictionary, the VQ mechanism acts as a robust information bottleneck. It effectively filters out phase-specific intensity variations and noise, ensuring that the extracted anatomical representation remains consistent regardless of whether the input is an arterial, venous, or non-contrast scan. We construct the learnable Embedding Dictionary $\mathcal{D}\in\mathbb{R}^{M\times C}$, containing $M$ anatomical vision queries, into which the anatomical information is distilled by quantization.
@@
where $k\in\{1,2,\dots,K-1\}$. This process reconstructs a refined multi-modal feature $\hat{X}_{s,N}$ by aggregating relevant queries from $\mathcal{D}$. To constrain the consistency of the anatomical information in multi-modal representations, we introduce pair-wise mean squared loss to the querying results:
@@ 3.6 Optimization Objectives
where $\mathcal{L}_{\text{DE}}$ is the disentanglement loss and $\mathcal{L}_{\text{Seg}}$ is the segmentation loss. Following [zhang2024prototypical], we use CLUB loss [cheng2020club] to disentangle the representations by mutual information (MI) minimization, which provides a contrastive upper bound on MI, preventing information leakage between the anatomical and temporal latent spaces. Specifically, we calculate
@@ 4.1 Datasets
The first dataset is the Climb 4 Kidney Cancer - Kidney and Kidney Tumor Segmentation (C4KC-KiTS) collection [C4KCKITS2019, heller2019kits19, heller2020state], a public subset of the KiTS19 challenge data. We utilized 210 patients from the training set. As shown in Fig. 4, every patient includes an arterial phase scan, while a subset of patients also contains non-contrast or delayed phase acquisitions, providing a controlled environment for testing missing-modality performance.
@@
While TARDis is theoretically grounded in the CT modalities, we extended our evaluation to the BraTS 2018 dataset [menze2014multimodal, bakas2017advancing, bakas2018identifying] to compare with existing state-of-the-art (SOTA) incomplete modality frameworks, which are predominantly established on this neuroimaging standard. BraTS18 provides four co-registered MRI modalities per patient: T1-weighted (T1), post-contrast T1-weighted (T1CE), T2-weighted (T2), and FLAIR. Annotations include three nested sub-regions: GD-enhancing tumor (ET), peritumoral edema (ED), and the necrotic/non-enhancing tumor core (NCR/NET). Following standard challenge protocols, we evaluated performance on three aggregated regions: Whole Tumor (WT), Tumor Core (TC), and Enhancing Tumor (ET).
@@ 4.3 Implementation Details
We compared TARDis against four leading methods in incomplete multi-modal learning: RFNet [ding2021rfnet], mmFormer [zhang2022mmformer], M3AE [liu2023m3ae], and M2FTrans [m2ftrans].
All experiments were conducted using PyTorch 1.12 on NVIDIA H20 GPUs. We utilized the AdamW optimizer with a decreasing learning rate schedule initialized at 0.01. The batch size was set to 4, employing the sequence-based optimization strategy detailed in Sec. 3.7. To simulate missing modalities during training, we applied a stochastic modality dropout with a rate of 20%, while ensuring that at least one modality remained available for every training sample. The embedding dictionary size was fixed at 512. The weight $\beta$ in Eq. 6 was set to 0.25, following [van2017neural]. When training with the BraTS18, the ranking loss $\mathcal{L}_{\text{Rank}}$ was replaced with a separation loss that attempts to push the $\tau$ far from each other. All baseline models were retrained and evaluated under identical conditions to ensure fair comparison.

### v3

# arXiv:2512.04576v3[cs.CV] 27 Feb 2026

@@ 1 Introduction
Tumor segmentation in contrast-enhanced Computed Tomography (CT) is fundamentally dependent on the transient physiological dynamics of the contrast agent [4]. Crucially, the diagnostic boundaries of tumors are often defined by this time-dependent contrast-enhancement signal, which is unavailable in non-contrast acquisitions or incorrectly timed scans [1]. Accurately capturing these dynamics usually requires a multi-phase acquisition protocol.
@@
However, obtaining a standardized, complete set of phases is rarely feasible in real-world clinical practice. First, strict limits on radiation dose and scanning time often force clinicians to skip specific phases, creating a significant missing modality problem. Second, and perhaps more challenging, is the variability of acquisition protocols across different medical centers and tumor types. The definition of a specific phase often relies on fixed time delays that do not account for patient-specific cardiac output or variations in scanner hardware. As a result, a scan labeled as arterial at one hospital may effectively be a late arterial or early venous phase at another. This temporal mismatch introduces severe domain shifts, where the intensity distribution of the same modality varies drastically between datasets. Standard deep learning frameworks, which assume fixed, synchronized input channels, degrade severely when facing such inconsistent or incomplete data [20].
@@
Existing approaches to multi-modal segmentation typically rely on channel-concatenation strategies, which assume a fixed set of inputs. When a modality is missing, these models often fail or require retraining. To address this, recent incomplete modality frameworks have been proposed, utilizing techniques such as arithmetic mean fusion, knowledge distillation, or shared latent space mapping [12, 7]. While these methods allow for flexible inputs, they generally treat different contrast phases as independent, uncorrelated channels or simple statistical variations. By ignoring the temporal dependency between phases, existing methods struggle to infer the missing dynamic information effectively, often resulting in suboptimal feature representation [28, 27].
@@
In this work, we propose Time Attenuated Representation Disentanglement (TARDis), a novel framework that addresses the incomplete modality challenge by modeling the underlying hemodynamics of CT. Our central hypothesis is that the Hounsfield Unit (HU) intensity can be explicitly decoupled into a time-invariant static component and a time-dependent dynamic component. Rather than treating missing modalities as empty data channels, we treat them as missing samples on the Time-Attenuation Curve (TAC). Our method employs a shared encoder to process a flexible number of input modalities. To learn robust representations, we introduce a dual-path disentanglement mechanism, as shown in Fig. 1. First, we extract a modal-agnostic anatomical representation using a learnable embedding dictionary, ensuring consistency across different scans. Second, we model the modal-specific dynamic representation by designing a physics-aware Conditional Variational Autoencoder (CVAE) [30], which is conditioned on the anatomy and the estimated scan time. This allows the network to probabilistically reconstruct the dynamic enhancement features of the tumor even when specific phases are missing. By explicitly disentangling anatomy from time-dependent perfusion, TARDis learns a physically grounded latent space that is robust to missing data, superior to standard fusion or masking strategies.
@@ 2.1 Multi-Modal Medical Image Segmentation
Multi-modal learning has become a cornerstone in medical image analysis, as combining complementary information from different modalities significantly improves diagnostic precision [35]. Early approaches focused on input-level fusion, where registered modalities are concatenated as multi-channel inputs to a Convolutional Neural Network [21, 25]. While effective, this assumes a fixed set of inputs, making the model fragile to missing data. Feature-level fusion strategies have subsequently emerged to better exploit cross-modal correlations. For instance, Dolz et al. [10] proposed HyperDenseNet to facilitate dense connectivity across multi-modal streams. More recently, Transformer-based architectures have been adopted to model long-range dependencies across modalities. Hatamizadeh et al. [13] introduced UNETR, which utilizes a vision transformer encoder to fuse volumetric data. Despite their success, these complete-modality frameworks rely on the assumption that all training and testing modalities are available. In clinical practice—particularly in multi-phase CT—scans are often incomplete due to radiation dose concerns or varying acquisition protocols, rendering these rigid fusion models ineffective without retraining or zero-filling strategies that often introduce artifacts.
@@ 2.2 Learning with Incomplete Modalities
To address the missing modality problem, recent research has shifted toward flexible frameworks capable of handling arbitrary subsets of inputs. The pioneering work of HeMIS [14] proposed learning a shared latent representation where first- and second-order statistics of available modalities are fused. This statistical fusion concept was extended by Dorent et al. [11] in the U-HVED framework, which integrates a multi-modal VAE to reconstruct missing inputs. Other approaches employ knowledge distillation (KD) to transfer information from a complete-modality teacher to an incomplete-modality student [18, 32]. However, KD methods typically require a complete set of modalities during the training phase, which may not always be available in retrospective cohorts. More recently, network architecture designs have moved toward adaptive feature fusion. RFNet [9] utilizes region-aware fusion to selectively aggregate features based on image content. With the rise of Transformers, methods like mmFormer [33] and M2FTrans [29] employ masked attention mechanisms to learn intra- and inter-modal correlations, effectively masking out missing tokens. M3AE [23] adapts the Masked Autoencoder (MAE) paradigm to multi-modal settings. While these methods achieve state-of-the-art results, they generally treat modalities as independent, uncorrelated channels.
@@ 2.3 Representation Disentanglement and Generative Modeling
Representation disentanglement aims to decompose data into distinct, explanatory factors, typically separating spatial structure from appearance. In natural image synthesis, methods like MUNIT [19] and DRIT++ [22] successfully disentangle content and style codes. In medical imaging, this paradigm has been applied to improve robustness and interpretability. Chartsias et al. [5] utilized disentanglement for synthesizing missing MRI sequences. SDNet [6] further decomposes medical images into spatial anatomy and non-spatial modality factors to improve segmentation generalization. VQ-VAE [31] demonstrates that quantizing latent features can effectively remove noise and preserve essential structural information. In the context of missing data, Generative Adversarial Networks and VAEs are often used to impute missing modalities before segmentation [36, 26]. However, most existing generative imputation methods perform image-to-image translation without explicit physical grounding.
@@ 3.3 Modal-Agnostic Representations
The static component represents the inherent anatomy, which should be consistent across all phases. However, features extracted from different contrast phases often contain phase-specific noise that obscures this underlying anatomy. We employ VQ-VAE [31] for the anatomical branch because anatomy is deterministic and time-invariant. By forcing features to map to a finite, discrete Embedding Dictionary, the VQ mechanism acts as a robust information bottleneck. It effectively filters out phase-specific intensity variations and noise, ensuring that the extracted anatomical representation remains consistent regardless of whether the input is an arterial, venous, or non-contrast scan. We construct the learnable Embedding Dictionary $\mathcal{D}\in\mathbb{R}^{M\times C}$, containing $M$ anatomical vision queries, into which the anatomical information is distilled by quantization.
@@
where $k\in\{0,1,\dots,K-1\}$. This process reconstructs a refined multi-modal feature $\hat{X}_{s,N}$ by aggregating relevant queries from $\mathcal{D}$. To constrain the consistency of the anatomical information in multi-modal representations, we introduce pair-wise mean squared loss to the querying results:
@@ 3.6 Optimization Objectives
where $\mathcal{L}_{\text{DE}}$ is the disentanglement loss and $\mathcal{L}_{\text{Seg}}$ is the segmentation loss. Following [34], we use CLUB loss [8] to disentangle the representations by mutual information (MI) minimization, which provides a contrastive upper bound on MI, preventing information leakage between the anatomical and temporal latent spaces. Specifically, we calculate
@@ 4.1 Datasets
The first dataset is the Climb 4 Kidney Cancer - Kidney and Kidney Tumor Segmentation (C4KC-KiTS) collection [16, 17, 15], a public subset of the KiTS19 challenge data. We utilized 210 patients from the training set. As shown in Fig. 4, every patient includes an arterial phase scan, while a subset of patients also contains non-contrast or delayed phase acquisitions, providing a controlled environment for testing missing-modality performance.
@@
While TARDis is theoretically grounded in the CT modalities, we extended our evaluation to the BraTS 2018 dataset [24, 2, 3] to compare with existing state-of-the-art (SOTA) incomplete modality frameworks, which are predominantly established on this neuroimaging standard. BraTS18 provides four co-registered MRI modalities per patient: T1-weighted (T1), post-contrast T1-weighted (T1CE), T2-weighted (T2), and FLAIR. Annotations include three nested sub-regions: GD-enhancing tumor (ET), peritumoral edema (ED), and the necrotic/non-enhancing tumor core (NCR/NET). Following standard challenge protocols, we evaluated performance on three aggregated regions: Whole Tumor (WT), Tumor Core (TC), and Enhancing Tumor (ET).
@@ 4.3 Implementation Details
We compared TARDis against four leading methods in incomplete multi-modal learning: RFNet [9], mmFormer [33], M3AE [23], and M2FTrans [29].
All experiments were conducted using PyTorch 1.12 on NVIDIA H20 GPUs. We utilized the AdamW optimizer with a decreasing learning rate schedule initialized at 0.01. The batch size was set to 4, employing the sequence-based optimization strategy detailed in Sec. 3.7. To simulate missing modalities during training, we applied a stochastic modality dropout with a rate of 20%, while ensuring that at least one modality remained available for every training sample. The embedding dictionary size was fixed at 512. The weight $\beta$ in Eq. 6 was set to 0.25, following [31]. When training with the BraTS18, the ranking loss $\mathcal{L}_{\text{Rank}}$ was replaced with a separation loss that attempts to push the $\tau$ far from each other. All baseline models were retrained and evaluated under identical conditions to ensure fair comparison.